Exercise 8-13 Direct Materials and Direct Labor Variances [LO8-4, LO8-5]

Huron Company produces a commercial cleaning compound known as Zoom. The direct materials and direct labor standards for one unit of Zoom are given below:


Standard Quantity or Hours Standard Price   
or Rate    
Standard
Cost   
  Direct materials 6.80 pounds $ 3.00 per pounds $ 20.40    
  Direct labor 0.50 hours $ 11.00 per hour $ 5.50    


During the most recent month, the following activity was recorded:
a. 22,900 pounds of material were purchased at a cost of $2.70 per pound.
b. All of the material purchased was used to produce 3,000 units of Zoom.
c. 1,400 hours of direct labor time were recorded at a total labor cost of $18,200.


Required:
1.

Compute the materials price and quantity variances for the month. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)

      

2.

Compute the labor rate and efficiency variances for the month. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Do not round your intermediate calculations.)
Answer #1
Standard Actual
Particulars Qty/ hours Rate amount Qty/ hours Rate amount
Materials              20,400.00                        3.00                 61,200.00              22,900.00                       2.70                 61,830.00
Labour                1,500.00                 11.0000                 16,500.00                1,400.00                     13.00                 18,200.00
Actual output                3,000.00
Materials reqd (3000 * 6.80)              20,400.00
Labour hrs reqd (3000*.50)                1,500.00
DMPV = (SP-AP)*AQ purchased
DMPV = (3 - 2.70)22,900
DMPV = 6,870 F
DMQV= (SQ-AQ)SP
DMQV= (20,400 - 22,900)3
DMQV= 7,500 U
DLRV= (SR-AR)AH
DLRV= (11 - 13)1,400
DLRV= 2,800 U
DLEV = (SH-AH)SR
DLEV = (1500 - 1400)11
DLEV = 1,100 F